Solve
5/3-2/x-4=1/2

Answer by rfer(16322)

-2/x=1/2+4-5/3
-2/x=3/6+24/6-10/6
-2x=17/6
x=(17/6)/-2
x=17/6*-1/2
x=-17/12
x=-1 5/12